40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 12 14:05:16 2014, ericq, Update, SUS, PRM Oplev Checked Out ol_spectra.pdf
    Reply  Wed Feb 12 14:59:36 2014, Steve, Update, SUS, clipping removed from PRM oplev PRMoplPointing.jpgPRMoplevReturn.jpg
       Reply  Wed Feb 12 19:37:05 2014, Jenne, Update, SUS, clipping removed from PRM oplev 
Message ID: 9627     Entry time: Wed Feb 12 14:05:16 2014     Reply to this: 9628
Author: ericq 
Type: Update 
Category: SUS 
Subject: PRM Oplev Checked Out 

 [ericq]

Steve fixed the PRM oplev pointing. I turned on the loops and measured the OLG, then set the pitch and yaw gains such that the upper UGF was ~8Hz (motivated by Jenne's loop design in ELOG 9401)

  • Pitch gain: +7
  • Yaw Gain: -5

I then measured the oplev spectra of the optics as they were aligned for PRMI. (OSEMs on, oplevs on, LSC off, and ASC off)

Next, Jenne and I need to fix the ASC loop such that it properly accounts for the oplev loop. 

ol_spectra.pdf 

 

ELOG V3.1.3-